### v1

# PersoBench: Benchmarking Personalized Response Generation in Large Language Models

Abstract
While large language models (LLMs) have exhibited impressive conversational capabilities, their proficiency in delivering personalized responses remains unclear. Although recent benchmarks automatically evaluate persona consistency in role-playing contexts using LLM-based judgment, the evaluation of personalization in response generation remains underexplored. To address this gap, we present a new benchmark, PersoBench, to evaluate the personalization ability of LLMs in persona-aware dialogue generation within a zero-shot setting. We assess the performance of three open-source and three closed-source LLMs using well-known datasets and a range of metrics. Our analysis, conducted on three well-known persona-aware datasets, evaluates multiple dimensions of response quality, including fluency, diversity, coherence, and personalization, across both standard and chain-of-thought prompting methods. Our findings reveal that while LLMs excel at generating fluent and diverse responses, they are far from satisfactory in delivering personalized and coherent responses considering both the conversation context and the provided personas. Our benchmark implementation is available at https://github.com/salehafzoon/PersoBench.
Keywords Personalized Response Generation  $\cdot$ Conversational AI, $\cdot$ Natural Language Processing $\cdot$ Large Language Models $\cdot$ Benchmarking
1 Introduction
Large Language Models (LLMs) have revolutionized NLP, excelling in human-like text generation across domains and becoming central to dialogue systems. However, evaluating their ability to generate personalized responses that enhance user engagement is crucial, especially in applications like customer service, where tailored interactions boost satisfaction [1].
While recent benchmarks such as RPBench-Auto [2], TIMECHARA [3] and RoleLLM [4] have been introduced in the role-playing domain to assess LLMs’ adherence to predefined characters or roles in character-based, scene-based, and temporal setups, there is still no dedicated benchmark for automatic personalized response generation of LLMs in the literature. Further, existing benchmarks also suffer from biases in their evaluations due to the use of large LLMs as judges, and limited experimental sizes constrain them.
To fill this gap, we introduce PersoBench, a benchmark for response personalization, to assess the strengths and limitations of current LLMs in generating personalized responses. To the best of our knowledge, no prior work has introduced a comprehensive benchmark specifically focused on evaluating response personalization in LLMs. Using comprehensive datasets and a diverse set of established metrics, including fluency, diversity, and coherence, we ensure a robust evaluation of various aspects of response generation, drawing on insights from a recent survey in the field [1]. More specifically, in line with this objective of the mentioned context, we aim to answer the following research questions:
  1. 1.
    Can LLMs generate fluent responses?
  2. 2.
    How proficient are LLMs at generating diverse responses?
  3. 3.
    Can LLMs generate responses that maintain coherence with the given context and persona?
  4. 4.
    Can LLMs generate personalized responses?
To answer the above questions, we benchmarked three open-source and three closed-source LLMs across three popular persona-aware conversational datasets. Their performance is evaluated in a zero-shot setting, considering both standard and Chain of Thought (COT) prompting configurations. Our evaluation goes beyond assessing overall response quality and contextualization, employing eight established metrics in conversational AI to gauge performance across fluency, diversity, coherence, and personalization dimensions. This work contributes by (1) offering a multi-aspect evaluation of LLM personalization in zero-shot settings, (2) investigating the influence of COT prompting on response personalization, and (3) assessing response time and instructability (the ability to adhere to task instructions) based on LLMs’ adherence to task instructions.
Our findings reveal that response personalization remains a challenging task for LLMs, highlighting the need for further research to enhance their performance in this regard.
2 Related Work
2.1 Existing Studies on LLM Benchmarking
LLM benchmarking has explored multiple aspects of their performance, including evaluations of factuality, ethical behaviour, task controllability, domain-specific applications, and cost efficiency. Besides their evaluation context, these benchmarks can be categorized based on the level of human involvement, which affects the degree of automation, and by their evaluation methods—whether they use explicit metrics for objective assessments or rely on LLMs as evaluators to judge model outputs.
Benchmarks relying on human-in-the-loop approaches require human involvement for tasks like error detection, ethical evaluation, and annotation. FELM [5], for example, uses human annotators to evaluate the factual accuracy of LLM outputs across domains like math, science, and reasoning. It focuses on granular annotations to capture subtle errors, providing detailed error types and reference links. TrustGPT [6] integrates human oversight for evaluating ethical concerns such as bias and value alignment, using tools like the PERSPECTIVE API [7] for toxicity scores while relying on human judgment for subjective issues.
On the other hand, some studies adopt the LLM-as-a-Judge approach for automated evaluation. Building on the need to assess LLMs in real-world tasks, DOCBench [8] benchmarks document reading systems from domains like academia, finance, and law using an LLM-based evaluation approach. The binary score assignment is done by GPT-4 model to evaluate the correctness of the executed task. While much of the question generation is automated, manual intervention is required for complex question types, limiting scalability. Expanding on specialized task benchmarking, PM-LLM-Benchmark [9] focuses on evaluating LLMs in process mining tasks, addressing the gap in assessing how well these models handle domain-specific knowledge. The benchmark evaluates open-source and commercial LLMs on tasks like process discovery, conformance checking, and root cause analysis. It uses two key implementation paradigms: direct insight provision and code generation. The evaluation relies on an LLM-as-a-Judge system, using GPT-4 to score outputs from 1 to 10. While the mentioned studies are effective and reduce human intervention, reliance on LLMs for evaluation can introduce bias, as LLMs may make subjective judgments.
Some studies have transitioned to automated evaluations by relying on explicit metrics instead of using LLMs as evaluators. CEBench [10] is an example of a fully automated evaluation toolkit, focusing on cost-effectiveness analysis of LLMs using metrics such as response time, memory efficiency, and mean absolute error (MAE). Its customizable setup helps optimise LLM pipelines for both effectiveness and budget and can be used to balance performance with financial considerations. Similarly, CoDI-Eval [11] explores the controllability of LLMs by assessing their ability to execute complex and varied instructions across key tasks such as sentiment, topic, length, keyword usage, and toxicity avoidance. The benchmark highlights the challenges LLMs face when adhering to explicit constraints and aims to improve alignment with human expectations. By integrating a broad range of instructions, CoDI-Eval evaluates both open-source and commercial models, including GPT-4 and Llama, using accuracy-based metrics to measure adherence to instructions without relying on subjective judgments or LLMs for scoring. These benchmarks demonstrate the potential to reduce human reliance while maintaining robust, measurable outcomes through explicit evaluation frameworks; however, they require specifically designed metrics tailored to different contexts for accurate assessments.
2.2 Existing Studies on Persona-aware Response Generation
Personalization, emerging as a crucial aspect of LLM usage, has been addressed through various evaluations, including published benchmarks. Focusing on released benchmarking, the release of RPBench-Auto by BosonAI [2] highlights this evaluation gap. The benchmark is built on persona-augmented conversations sourced from online communities. Inspired by ArenaHard [12] and AlpacaEval [13], RPBench-Auto uses GPT-4-Turbo as a judge model to compare responses from evaluated models against a GPT-4o baseline in character-based and scene-based role-plays. Although effective, this approach is prone to biases inherent in both the baseline and judge models. Their evaluation relies on a win rate metric over a limited test set (80 personae), rather than explicit evaluation metrics, to measure how often the evaluated model’s responses are preferred over the baseline. The study underscores the pressing need for a specialized benchmark in personalized response generation, and an analysis of the CoT effect could further enhance their setup. In the same context, TIMECHARA [3] benchmarks the temporal consistency of character personas in narratives and the potential hallucinations that can occur in LLMs, focusing on how well these models maintain character integrity over time. The evaluation framework introduces metrics like AlignScore to assess spatiotemporal consistency and a 1-7 Likert scale [14] to measure personality consistency. These evaluations are conducted across various setups, including zero-shot, few-shot, CoT reasoning, and retrieval-augmented generation (RAG). However, the use of LLMs as judges introduces potential biases and increases costs, limiting the feasibility of large-scale assessments. In a related effort, RoleLLM [4] explores a more fine-tuned role-conditioning process to benchmark models’ ability to imitate characters with specific traits. Both open-source and closed-source models are evaluated using a custom dataset, RoleBench. The evaluation incorporates Rouge-L for role-playing accuracy and GPT-4-based human-like evaluations to rank models on their persona-specific performance. While effective at eliciting character-level responses, RoleLLM faces limitations such as reliance on single-turn interactions and manual intervention in data generation, which may affect scalability and the richness of dialogue.
Figure 1: Overview of the PersoBench automatic personalization benchmark framework.
In addition to benchmarking, recent evaluation studies using datasets such as IT-ConvAI2, Blended Skill Talk, and FoCus—also integrated into our work—have focused on assessing dialogue models. For example, the research by Chen et al. (2023) [15] focuses on metrics like BLEU and ROUGE, which, while informative, fail to fully capture the personalization capabilities of more advanced models. Innovative approaches, such as PAL [16] and robust personalized dialogue generation [17], have pushed the field forward, yet the need for more comprehensive evaluation metrics remains evident.
However, the mentioned studies often fall short due to their reliance on a limited array of metrics and older models like GPT-2. Additionally, these studies are constrained by their narrow scope and evaluation frameworks. In contrast, our work comprehensively evaluates both open-source and closed-source LLMs, employing dedicated metrics to assess personalization, fluency, diversity, and coherence. By analyzing model performance in both vanilla and CoT setups within a zero-shot setting, we provide an in-depth comparison across these aspects, setting a new benchmark for evaluating response personalization in LLMs.
3 PersoBench
3.1 Problem statement
For our benchmark on personalized response generation, we have adopted the same problem statement as defined by Chen et al. (2023) [17]. Given the dialogue context $C=\{u_{1},\dots,u_{m}\}$ and a set of persona descriptions $P=\{p_{1},\dots,p_{n}\}$, the goal is to generate a personalized response $r$. Formally, the generation problem can be formulated as the following chain rule:
$P(r\mid C,P;\theta)=\prod_{t=1}^{T}P(r_{t}\mid r_{1:t-1},C,P;\theta)$ (1)
where $\theta$ is the parameters of the dialogue model.
3.2 Overview of PersonBench
Our automatic benchmarking framework is shown in Fig.1. PersoBench includes around 3,600 samples from three persona-aware datasets, each differing in size and context. It applies eight metrics across four evaluation dimensions based on the datasets and metrics discussed in the recent review study [1]. The framework evaluates the performance of six prominent LLMs—three open-source and three closed-source— within a zero-shot setting under both vanilla and COT prompting setups.
3.3 Prompt Development
In developing prompts aligned with implicit design principles and representative examples from the literature, we adhered to commonly accepted practices, although no explicit standards exist. To preserve the integrity of our benchmarking study, we intentionally refrained from employing advanced prompt engineering techniques that could obscure common challenges faced by LLMs, such as hallucinations, inaccuracies, and generalizations.
Prompts Structure
I will provide you with a conversation context and the personas of the participants, that can be annotated with speaker information. As a participant in this conversation, your task is to generate a personalized response, considering the conversation context and the annotated personas. Participant Personas: {personas} Conversation Context: {conversation history} Task Instruction: * Provide an unannotated response. * If only one persona is available, personalize the response accordingly. * If the conversation context is a single query, respond appropriately to the query. * Apply Chain of Thought reasoning to reflect on the alignment of your response with the personas. Output Format: Output a JSON of the following format: { "reasoning": "briefly describe your personalization process (in 110 words or less)." "response": "provide the personalized natural language response here." }
Table 1: Blueprint of the designed prompt for personalized response generation task. The orange text will only be included in the COT setup.
The overall structure of our designed prompt is shown in Table 1. As suggested by related studies [18, 19, 20, 5], the prompt begins with a general problem description to establish context for the LLM, followed by a clear specification of the LLM’s role in addressing the target task. Next, a clear definition of the input is provided, which includes participants’ personas and conversation history, ensuring the necessary information is available for the task. In line with the principles outlined in the literature [21, 22, 18, 23], detailed task instructions are given to guide response generation and mitigate ambiguities. Lastly, the prompt specifies the desired output format for the LLM, supporting both vanilla and COT setups. Specifying the JSON format is crucial for our benchmarking setup, allowing seamless extraction of reasoning and responses for post-processing and evaluation. This strategy is also commonly applied when the output requires more than a simple response [24, 25]. In scenarios where the conversation context involves IT-related queries, our experiments have shown that instructing the LLM to provide a ’natural language response’ is crucial. This prevents the model from generating code instead of a straightforward conversational answer. Based on findings that placing the reasoning before the response improves coherence and adherence to task requirements [24], we adopted this approach in the design of the output structure. To ensure the reasoning does not consume all available tokens and leave insufficient space for the actual response, we included the keywords ’briefly’ and ’(in less than 110)’ in the reasoning instruction. The 110-token limit was determined based on the longest actual response across all selected datasets (shown in Table 2), ensuring an equal token allocation for both the reasoning and the response.
4 Experiment
4.1 Dataset
To establish a solid foundation for our benchmark, we selected widely recognized datasets in the domain, guided by a recent survey in the same context [1]. Our selection criteria included choosing datasets with a reasonable test/validation set size, which was important given our budget constraints. We prioritized datasets that clearly separated the actual response from the conversation context to meet our data structure requirements and those that required minimal preprocessing to preserve the integrity of their content. Based on these criteria, the following datasets were chosen for our benchmarking experiment, and their statistics are provided in Table 2.
Blended Skill Talk (BST): The BST dataset [^1] https://www.kaggle.com/datasets/thedevastator/multi-modal-conversation-data?resource=download evaluates key conversational competencies, including dialogue flow, topic control, coherence, and the integration of personality, empathy, and knowledge. This dataset is effective for assessing conversational styles, enhancing user engagement, and benchmarking distributed dialogue systems across various modalities.
Follow-up Customized conversation (FoCus) [^2] https://github.com/pkchat-focus/FoCus?tab=readme-ov-file This dataset is designed to generate personalized and knowledgeable responses by grounding them in both the user’s persona and relevant Wikipedia knowledge. Focused on conversations about geographical landmarks, it evaluates how well models integrate persona and knowledge to create customized, contextually appropriate responses, making it valuable for developing advanced conversational agents.
IT-ConvAI2: [^3] https://github.com/CCIIPLab/Persona_Extend/tree/main/datasets/IT_ConvAI2 As an extension of the original ConvAI2 dataset, it is tailored specifically for IT-related conversations. It focuses on dialogues in technical support scenarios, where the conversational agent is expected to assist users with IT issues. As a persona-driven dataset, it includes personalized information about the participants to make the interactions more relevant and context-aware.
Table 2: Characteristics of the selected datasets for benchmarking. ’Resp. Len.’ indicates the overall golden response length range in words.
Name Test Size Context Type Query Sep U/D Role Resp. Len.
BST 1k Dialogue Yes 11.6 H-H 1-30
FoCus 1k Dialogue No 11.3 H-A 4-108
IT-ConvAI2 1.6k Query - 2.0 H-H 4-22
Table 3: Fluency evaluation of LLMs on the IT-ConvAI2 dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.12/0.02/0.11 0.14 0.81/0.81/0.81 0.09/0.01/0.09 0.11 0.80/0.80/0.80
Qwen2 7B 0.10/0.02/0.10 0.14 0.83/0.84/0.83 0.09/0.01/0.08 0.12 0.82/0.83/0.83
Llama3.1 8B 0.11/0.02/0.10 0.13 0.76/0.77/0.76 0.11/0.02/0.10 0.12 0.84/0.83/0.84
GPT3.5 Turbo 0.12/0.02/0.11 0.15 0.85/0.85/0.85 0.11/0.02/0.10 0.13 0.84/0.84/0.84
GPT4o Mini 0.11/0.02/0.10 0.16 0.84/0.85/0.84 0.10/0.01/0.10 0.16 0.83/0.85/0.84
GPT4 Turbo 0.12/0.02/0.11 0.15 0.85/0.86/0.85 0.11/0.02/0.10 0.13 0.84/0.84/0.84
Table 4: Fluency evaluation of LLMs on the Blended Skill Talk dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.10/0.01/0.10 0.11 0.80/0.81/0.81 0.10/0.01/0.09 0.11 0.80/0.82/0.81
Qwen2 7B 0.11/0.01/0.10 0.12 0.82/0.84/0.83 0.09/0.01/0.09 0.11 0.83/0.86/0.84
Llama3.1 8B 0.12/0.02/0.11 0.12 0.81/0.82/0.82 0.11/0.01/0.10 0.11 0.85/0.86/0.85
GPT3.5 Turbo 0.13/0.02/0.11 0.13 0.85/0.86/0.86 0.11/0.01/0.10 0.12 0.82/0.84/0.83
GPT4o Mini 0.13/0.02/0.12 0.14 0.85/0.87/0.86 0.12/0.01/0.11 0.14 0.84/0.86/0.85
GPT4 Turbo 0.12/0.02/0.11 0.12 0.85/0.86/0.86 0.11/0.01/0.10 0.12 0.84/0.86/0.85
4.2 Implementation setup
Pre-Processing: To ensure consistency and reliability in our benchmarking process, we established a basic pre-processing workflow across all datasets. This workflow included simple text cleaning. Notably, we did not alter the inherent structure of the data, as the separation between persona and context was already well-defined across all the selected datasets. We added missing context annotations to ensure the LLM can distinguish speaker turns, which is essential for clarity. For persona annotations, we retained the original format provided in the datasets. In cases where only one persona is given, this does not impact the LLM’s ability to personalize responses, as it is instructed to do so based on the provided persona.
Post-processing Instructing LLMs to generate output in JSON format (Table 1), any unparsable outputs are considered failed instances and factored into the failure ratio calculation. From the successfully parsed outputs, we extracted only the generated response to compare against the golden response, discarding the reasoning portion, as the analysis focused on the impact of COT reasoning on the response itself. Additionally, response generation time, measured in seconds, provides valuable insights for practical industrial applications, particularly when deploying open-source LLMs.
Metrics: We evaluate the generated responses across fluency, diversity, coherence, and personalization. Fluency is measured using ROUGE [26], METEOR [27], and BERTScore [28] to assess linguistic quality and semantic similarity. Diversity is evaluated using Dist-1 and Dist-2 [29] to calculate the ratio of unique unigrams and bigrams. For personalization, the Consistency Score (C Score) [30] measures persona consistency, while the Persona Distance (P-Dist) score [31] assesses persona coverage. Coherence is evaluated with the Utterance Entailment Score (UE-Score) [32] based on Natural Language Inference (NLI). These metrics ensure a comprehensive evaluation of the model’s performance.
LLMS: To implement an automatic evaluation pipeline, we chose the instructed versions of LLMs. This allows us to automate content extraction from the generated responses, streamlining the post-processing and evaluation modules. In this regard, we selected Mistral 7B [^4] https://mistral.ai/news/announcing-mistral-7b/, Qwen2 7B [^5] https://github.com/QwenLM/Qwen2, and Llamma3.1 8B [^6] https://llama.meta.com/ as the three open-source LLMs. For the closed-source models, we utilized the GPT series from OpenAI [^7] https://openai.com/research/gpt, including GPT-3.5 Turbo, GPT-4 Turbo, and GPT-4o Mini.
4.3 Expriment Results
Table 5: Fluency evaluation of LLMs on the FoCus dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.11/0.03/0.10 0.11 0.48/0.49/0.48 0.15/0.03/0.13 0.15 0.75/0.77/0.76
Qwen2 7B 0.08/0.02/0.08 0.10 0.41/0.43/0.42 0.10/0.02/0.09 0.13 0.60/0.63/0.62
Llama3.1 8B 0.11/0.03/0.10 0.11 0.47/0.47/0.47 0.15/0.03/0.13 0.16 0.80/0.81/0.80
GPT3.5 Turbo 0.20/0.06/0.18 0.20 0.78/0.79/0.78 0.17/0.04/0.15 0.18 0.79/0.80/0.80
GPT4o Mini 0.19/0.06/0.17 0.22 0.78/0.81/0.80 0.20/0.05/0.17 0.22 0.85/0.87/0.86
GPT4 Turbo 0.23/0.08/0.20 0.23 0.84/0.84/0.84 0.22/0.07/0.19 0.23 0.85/0.86/0.86
Table 6: Diversity evaluation of LLMs using Dist-1 and Dist-2 metrics across benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2
Mistral 7B 0.86 0.95 0.87 0.95 0.87 0.94 0.86 0.95 0.49 0.56 0.75 0.88
Qwen2 7B 0.90 0.98 0.90 0.99 0.90 0.97 0.91 0.99 0.43 0.49 0.60 0.72
Llama3.1 8B 0.78 0.89 0.89 0.99 0.84 0.95 0.90 0.99 0.44 0.53 0.76 0.93
GPT3.5 Turbo 0.91 1.00 0.91 0.99 0.92 1.00 0.87 0.97 0.79 0.90 0.78 0.92
GPT4o Mini 0.89 0.99 0.88 0.99 0.91 1.00 0.89 0.99 0.79 0.91 0.85 0.99
GPT4 Turbo 0.93 1.00 0.93 1.00 0.92 1.00 0.91 1.00 0.86 0.96 0.86 0.98
Can LLMs generate fluent responses? The benchmarked LLMs generally demonstrate strong fluency across various datasets, regardless of context length. This is particularly evident in Tables 3, 4 and 5, in the high BERTScore values, which show that the models effectively capture the meaning and structure of conversations. While ROUGE and METEOR scores indicate less reliance on exact phrase matches with the golden responses, this does not detract from overall fluency. BERTScore proves to be a more practical metric in conversational AI, where capturing the intended meaning is more important than exact word matches.
Both groups perform well when comparing open-source and closed-source LLMs, with closed-source models like those in the GPT-4 series slightly outperforming their open-source counterparts, particularly in more extended contexts such as those in the FoCus dataset. However, the performance gap is not substantial, showing that open-source models are still capable, especially in less complex dialogues.
Among the open-source LLMs, Qwen2 performs well in shorter contexts, while Llama3.1 excels in longer and more challenging dialogues, such as the FoCus dataset. Llama3.1’s larger parameter size and effective use of COT reasoning allow it to handle more complex contexts, surpassing other open-source models. Meanwhile, GPT-4 models consistently demonstrate robust fluency across all contexts. Although COT generally enhances their fluency in shorter contexts, it can sometimes introduce unnecessary complexity, slightly reducing performance by making responses more rigid and formulaic. Despite this, GPT-4 models maintain strong overall fluency in both short and long dialogues.
How proficient are LLMs at generating diverse responses? Benchmarked LLMs are generally good at generating diverse responses, particularly in shorter-context datasets like IT-ConvAI2 and Blended Skill Talk (Table 6). This is supported by high Dist-1 and Dist-2 scores, which measure the diversity of unigrams and bigrams, indicating a wide variety of words and varied word combinations in the generated responses. When comparing the closed-source and open-source LLMs, closed-source models like those in the GPT-4 series tend to perform slightly better overall, especially in longer contexts. However, the difference between the two groups is not stark, indicating that open-source models are still quite capable, particularly in less complex dialogues.
For open-source LLMs, performance tends to decline as the context length increases, as seen in the FoCus dataset, which involves longer and more complex dialogues. This suggests that while these models handle diversity well in shorter exchanges, they face challenges maintaining the same diversity level in longer contexts.
The impact of COT is particularly noticeable in open-source LLMs when dealing with longer contexts. For instance, Llama3.1 benefits the most from COT, likely due to its larger parameter size, which helps manage more intricate dialogues, as it has an extra billion parameters compared to other open-source LLMs. On the other hand, closed-source LLMs like those from the GPT-4 series show minimal improvement with COT. In some cases, especially in shorter contexts, their performance even slightly degrades. This may occur because COT can cause these models to overstructure their responses, leading to less natural diversity and producing too rigid or formulaic outputs. Despite this, the GPT-4 generation consistently stands out as top performers across varying context lengths, demonstrating robust diversity generation in both short and long dialogues.
Table 7: Coherence evaluation using UE Score on benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Mistral 7B 0.46 0.31 0.48 0.44 0.21 0.33
Qwen2 7B 0.42 0.38 0.39 0.52 0.17 0.36
Llama3.1 8B 0.23 0.22 0.33 0.37 0.17 0.33
GPT 3.5 Turbo 0.49 0.38 0.51 0.57 0.37 0.45
GPT 4o Mini 0.40 0.38 0.38 0.39 0.35 0.37
GPT4 Turbo 0.43 0.42 0.56 0.50 0.39 0.45
Table 8: Personalization evaluation of LLMs using P-Dist Score and C Score across the benchmarking dataset, in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score P-Dist Score C Score
Mistral 7B 0.62 0.33 0.54 0.02 0.46 -0.00 0.43 -0.08 0.30 -0.23 0.53 0.41
Qwen2 7B 0.61 0.21 0.63 0.23 0.46 -0.11 0.49 0.01 0.29 -0.31 0.43 0.21
Llama3.1 8B 0.55 0.03 0.58 0.05 0.45 -0.11 0.46 -0.05 0.33 -0.22 0.59 0.54
GPT3.5 Turbo 0.63 0.23 0.60 0.10 0.47 -0.01 0.48 0.02 0.51 0.26 0.52 0.40
GPT4o Mini 0.63 0.10 0.64 0.14 0.48 -0.10 0.50 -0.05 0.54 0.16 0.57 0.30
GPT4 Turbo 0.60 0.08 0.59 0.01 0.47 -0.05 0.47 -0.05 0.51 0.14 0.52 0.21
Can LLMs generate responses that maintain coherence with the given context and persona? Based on the UE score presented in Table 7, which is calculated using an NLI model whic relies on natural language inference, LLMs generally struggle with generating coherent responses, especially when working with long-context datasets like FoCus. Even in shorter-context datasets, as seen in the IT-ConvAI2 results, LLMs show poor performance in maintaining coherence. While the closed-source LLMs perform better at this task, even powerful models like GPT-4-Turbo fall short of achieving good results. Among them, GPT4o Mini demonstrates marginally lower performance across all the datasets with varying context sizes. While COT reasoning offers a modest performance improvement, GPT series LLMs, as fine-tuned LLMs, remain limited in their ability to handle the coherence task effectively.
Although the open-source LLMs’ performances are closely aligned, Llama3.1 exhibits marginally lower results across the datasets, indicating its inability to leverage the additional billion parameters effectively. It also benefits less from COT compared to the others. Overall, open-source LLMs face significant challenges in maintaining coherence, particularly in complex, long-context scenarios.
Can LLMs generate personalized responses? The results of P-Dist and C scores, which are shown in Table 8, two well-established NLI-based metrics, reveal that LLMs struggle significantly with personalization. There is no clear benefit in using closed-source LLMs over open-source alternatives, underscoring the inherent difficulty of the personalization task for these models. Even in the best scenarios, the P-Dist score—where 1 represents the ideal outcome—demonstrates only moderate performance. The C score, which also penalizes contradictory entailments with negative values, exposes these limitations even more starkly, particularly in short-length contexts where less content is available for personalization. On short-context datasets, no clear performance trend emerges among open-source LLMs. However, on the FoCus dataset, Llama3.1 shows a slight superiority in the personalization task. The same overall scenario applies for close-source LLMs, while GPT-4 Turbo emerges as the least effective performer. COT negatively impacts short-length conversations but yields a notable improvement in long-length conversations, as seen with the FoCus dataset. This suggests that the added complexity of COT is only advantageous when there is enough content to work with. This need for COT reasoning aligns with findings from [33], which highlight that current language models do not reliably utilize information in long input contexts.
5 Analysis and Discussion
To further analyse the LLM performance on longer text, we focused on the FoCus dataset. Our previous results show that its long-context nature makes LLM performance differences more apparent. Fig. 2 provides an overall view of LLM groups across different prompting configurations. BertScore-F1, Dist-2, UE Score, P-Dist, and C Score are considered for overall fluency, diversity, coherence, persona coverage and persona consistency analysis. Response personalization remains challenging, even for fine-tuned closed-source LLMs, as evidenced by the results in both persona consistency and persona coverage. Additionally, the coherence evaluation, which accounts for both persona and golden response alignment, further highlights this difficulty. While open-source LLMs generally cannot compete with closed-source models in terms of fluency and diversity, their performance in personalization and coherence is considerably enhanced with the use of COT. In fact, an open-source model like Llama3.1 can even surpass fine-tuned closed-source LLMs in these areas. Meanwhile, the performance of closed-source LLMs do not show any tangible improvement.
((a))
((b))
((c))
((d))
Figure 2: Performance analysis of (a) open-source LLMs in vanilla setting (b) close-source LLMs in vanilla setting (c) open-source LLMs in COT setting (b) close-source LLMs in COT setting on the FoCus dataset.
((a))
((b))
Figure 3: Analysis of response time and failure ratio for LLMs on the FoCus dataset: (a) Vanilla setup and (b) COT setup.
Fig.3 illustrates LLMs’ comparative performance from response time and failure ratio, two aspects that can provide information for practical insights for selecting LLMs for industrial use. The failure ratio can indicate the model’s instructability, which refers to how well the model follows instructions, particularly considering the maximum token limit provided during response generation. While COT increases response time, it directly enhances the instructability of both open-source and closed-source LLMs, with a significant difference in their ability to handle varying token length limits.
6 CONCLUSION AND FUTURE WORK
In this paper, we introduced PersoBench, an automatic benchmarking pipeline designed to evaluate LLMs in the context of persona-aware conversation generation. Using both vanilla and COT prompting setups in a zero-shot setting, we benchmarked well-known open-source and closed-source LLMs on three persona-aware datasets. Leveraging eight well-established evaluation metrics, we assessed the models across fluency, diversity, coherence, and personalization dimensions. Our findings suggest that while current LLMs perform well in generating fluent and diverse responses, they struggle to deliver personalized and coherent responses in persona-augmented conversational contexts.
A potential future direction for research in this domain could involve exploring beyond textual persona representation to include other forms, such as tabular demographic data or multi-modal setups incorporating different media types.
Limitations
While significant effort has been invested in this research, there are still some limitations to our study: (1) In crafting an automatic evaluation pipeline, we required the LLMs to format their generated output in JSON format to enable the automatic extraction of the response and reasoning. However, this automation discards responses that could not be successfully parsed as JSON, necessitating further manual intervention and potentially extending the post-processing workload with human resources. (2) To increase the likelihood of successful response generation in JSON format, we opted for instruct-tuned versions of LLMs. While this choice was necessary for the study, it excludes the evaluation of standard versions of LLMs that may perform well in more general, unstructured conversational contexts.
Acknowledgments
We acknowledge the Centre for Applied Artificial Intelligence at Macquarie University (Sydney, Australia) for funding this research.
A Better LLM Evaluator for Text Generation[24] Data: Here are some conversations that happened on the same day: [Fill in the conversations]
Task: Please evaluate the conversations considering three aspects: factual inconsistencies, redundancies, and illogical statements.
Scoring: Assign an evaluation score on a scale from 1 to 10, where:
* 1 indicates that there are no issues.
* 10 suggests the presence of numerous issues, particularly in the aspects of factual inconsistencies, redundancies, and illogical statements.
Special Rules: Do not give a small score just because the issue is not very impactful. Consider the number of issues rather than its impact. Your overall score should reflect the utmost concern observed in any of the aspects.
Output Format: Output a JSON of the following format:
"reasons": "point out the issues and your reasons for the rating",
"score": "<json integer>"
Table 9: Explicit Prompt Examples.
Sample Prompt
I will provide you with a conversation context and the personas of the participants, that can be annotated with speaker information. As a participant in this conversation, your task is to generate a personalized response, considering the conversation context and the annotated personas. Participant Personas: [User 1 persona]: [’i attend book club every week.’ ’i am more of a cat person than a dog person.’] Conversation Context: User1: my parents dies in an accident User2: that is sad . my best friend is my mom. i could not image loosing her. User1: it is harder some days than others. but it is good you get along with your mom so well. Task Instruction: * Provide an unannotated response. * If only one persona is available, personalize the response accordingly. * If the conversation context is a single query, respond appropriately to the query. * Apply Chain of Thought reasoning to reflect on the alignment of your response with the personas. Output Format: Output a JSON of the following format: { "reasoning": "briefly describe your personalization process (in 110 words or less)." "response": "provide the personalized natural language response here." }
Table 10: Sample COT setup prompt.
7.2 Metrics
To evaluate fluency, we used ROUGE [26], METEOR [27], and BERTScore [28]. ROUGE, via the rouge-score package [^8] ROUGE-Score package: https://github.com/google-research/google-research/tree/master/rouge, measures n-gram overlap between generated and reference texts. METEOR, implemented through NLTK [^9] NLTK package: https://www.nltk.org/, accounts for synonymy and stemming, offering a more nuanced linguistic evaluation. BERTScore, implemented using the bert-score package [^10] BERTScore package: https://github.com/Tiiiger/bert_score, evaluates Precision, Recall, and F1 based on contextual embeddings, assessing how well the generated text preserves meaning and coherence relative to the reference.
For diversity assessment, we calculated Dist-1 and Dist-2 metrics [29], which respectively measure the ratio of unique unigrams and bigrams in the generated outputs.
Additionally, We used advanced metrics to evaluate personalization and coherence in conversational agents. Personalization can be evaluated from two aspects: consistency and coverage. Consistency reflects whether the generated response R is aligned with the given personal information P, while coverage indicates how much of the given personal information is reflected in the generated response. The C Score [30], a metric for persona consistency, evaluates whether the responses remain consistent with the persona by fine-tuning BERT on Dialogue Natural Language Inference (DNLI).
On the other hand, the Persona Distance [31], a metric for persona coverage, measures the semantic alignment between the response and persona by utilizing word2vec cosine similarity. Additionally, the Utterance Entailment Score (UE-Score) [32] measures dialogue coherence by calculating the NLI alignment score between an utterance and its response. The UE-Score model is based on BERT, which has been fine-tuned using the Stanford Natural Language Inference (SNLI) [^11] https://nlp.stanford.edu/projects/snli/ dataset [34]. These metrics are essential for a nuanced assessment of personalized dialogue systems.
Table 11: Fluency evaluation of LLMs on the IT-ConvAI2 dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.12 $\pm$ 0.1 / 0.02 $\pm$ 0.05 / 0.11 $\pm$ 0.09 0.14 $\pm$ 0.11 0.81 $\pm$ 0.18 / 0.81 $\pm$ 0.18 / 0.81 $\pm$ 0.18 0.09 $\pm$ 0.08 / 0.01 $\pm$ 0.04 / 0.09 $\pm$ 0.07 0.11 $\pm$ 0.09 0.80 $\pm$ 0.18 / 0.80 $\pm$ 0.18 / 0.80 $\pm$ 0.18
Qwen2 7B 0.10 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.08 0.14 $\pm$ 0.1 0.83 $\pm$ 0.12 / 0.84 $\pm$ 0.12 / 0.83 $\pm$ 0.12 0.09 $\pm$ 0.06 / 0.01 $\pm$ 0.03 / 0.08 $\pm$ 0.06 0.12 $\pm$ 0.09 0.82 $\pm$ 0.08 / 0.83 $\pm$ 0.08 / 0.83 $\pm$ 0.08
Llama3.1 8B 0.11 $\pm$ 0.09 / 0.02 $\pm$ 0.05 / 0.10 $\pm$ 0.09 0.13 $\pm$ 0.10 0.76 $\pm$ 0.25 / 0.77 $\pm$ 0.25 / 0.76 $\pm$ 0.25 0.11 $\pm$ 0.09 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.08 0.12 $\pm$ 0.09 0.84 $\pm$ 0.07 / 0.83 $\pm$ 0.07 / 0.84 $\pm$ 0.07
GPT3.5 Turbo 0.12 $\pm$ 0.09 / 0.02 $\pm$ 0.05 / 0.11 $\pm$ 0.08 0.15 $\pm$ 0.11 0.85 $\pm$ 0.02 / 0.85 $\pm$ 0.03/ 0.85 $\pm$ 0.02 0.11 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.07 0.13 $\pm$ 0.10 0.84 $\pm$ 0.03 / 0.84 $\pm$ 0.03 / 0.84 $\pm$ 0.03
GPT4o Mini 0.11 $\pm$ 0.07 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.07 0.16 $\pm$ 0.10 0.84 $\pm$ 0.05 / 0.85 $\pm$ 0.05 / 0.84 $\pm$ 0.05 0.10 $\pm$ 0.06 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.06 0.16 $\pm$ 0.09 0.83 $\pm$ 0.04 / 0.85 $\pm$ 0.04 / 0.84 $\pm$ 0.04
GPT4 Turbo 0.12 $\pm$ 0.09 / 0.02 $\pm$ 0.05 / 0.11 $\pm$ 0.08 0.15 $\pm$ 0.11 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.03 / 0.85 $\pm$ 0.02 0.11 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.08 0.13 $\pm$ 0.10 0.84 $\pm$ 0.02 / 0.84 $\pm$ 0.02 / 0.84 $\pm$ 0.02
Table 12: Fluency evaluation of LLMs on the Blended Skill Talk dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.10 $\pm$ 0.08 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.11 $\pm$ 0.08 0.80 $\pm$ 0.19 / 0.81 $\pm$ 0.19 / 0.81 $\pm$ 0.19 0.10 $\pm$ 0.07 / 0.01 $\pm$ 0.02 / 0.09 $\pm$ 0.07 0.11 $\pm$ 0.08 0.80 $\pm$ 0.18 / 0.82 $\pm$ 0.18 / 0.81 $\pm$ 0.18
Qwen2 7B 0.11 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.12 $\pm$ 0.08 0.82 $\pm$ 0.13 / 0.84 $\pm$ 0.14 / 0.83 $\pm$ 0.13 0.09 $\pm$ 0.06 / 0.01 $\pm$ 0.02 / 0.09 $\pm$ 0.06 0.11 $\pm$ 0.07 0.83 $\pm$ 0.05 / 0.86 $\pm$ 0.05 / 0.84 $\pm$ 0.05
Llama3.1 8B 0.12 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.11 $\pm$ 0.08 0.12 $\pm$ 0.09 0.81 $\pm$ 0.17 / 0.82 $\pm$ 0.18 / 0.82 $\pm$ 0.17 0.11 $\pm$ 0.08 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.11 $\pm$ 0.08 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.85 $\pm$ 0.02
GPT3.5 Turbo 0.13 $\pm$ 0.09 / 0.02 $\pm$ 0.04 / 0.11 $\pm$ 0.08 0.13 $\pm$ 0.09 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.86 $\pm$ 0.02 0.11 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.06 0.12 $\pm$ 0.08 0.82 $\pm$ 0.13 / 0.84 $\pm$ 0.13 / 0.83 $\pm$ 0.13
GPT4o Mini 0.13 $\pm$ 0.08 / 0.02 $\pm$ 0.03 / 0.12 $\pm$ 0.07 0.14 $\pm$ 0.09 0.85 $\pm$ 0.03 / 0.87 $\pm$ 0.04 / 0.86 $\pm$ 0.03 0.12 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.11 $\pm$ 0.06 0.14 $\pm$ 0.08 0.84 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.85 $\pm$ 0.02
GPT4 Turbo 0.12 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.11 $\pm$ 0.07 0.12 $\pm$ 0.09 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.86 $\pm$ 0.02 0.11 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.12 $\pm$ 0.08 0.84 $\pm$ 0.03 / 0.86 $\pm$ 0.04 / 0.85 $\pm$ 0.03
Table 13: Fluency evaluation of LLMs on the FoCus dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.11 $\pm$ 0.12 / 0.03 $\pm$ 0.06 / 0.10 $\pm$ 0.11 0.11 $\pm$ 0.14 0.48 $\pm$ 0.42 / 0.49 $\pm$ 0.43 / 0.48 $\pm$ 0.42 0.15 $\pm$ 0.09 / 0.03 $\pm$ 0.04 / 0.13 $\pm$ 0.08 0.15 $\pm$ 0.10 0.75 $\pm$ 0.25 / 0.77 $\pm$ 0.26 / 0.76 $\pm$ 0.26
Qwen2 7B 0.08 $\pm$ 0.10 / 0.02 $\pm$ 0.04 / 0.08 $\pm$ 0.09 0.10 $\pm$ 0.13 0.41 $\pm$ 0.42 / 0.43 $\pm$ 0.44 / 0.42 $\pm$ 0.43 0.10 $\pm$ 0.08 / 0.02 $\pm$ 0.03 / 0.09 $\pm$ 0.07 0.13 $\pm$ 0.10 0.60 $\pm$ 0.37 / 0.63 $\pm$ 0.38 / 0.62 $\pm$ 0.38
Llama3.1 8B 0.11 $\pm$ 0.15 / 0.03 $\pm$ 0.1 / 0.10 $\pm$ 0.14 0.11 $\pm$ 0.14 0.47 $\pm$ 0.42 / 0.47 $\pm$ 0.43 / 0.47 $\pm$ 0.43 0.15 $\pm$ 0.10 / 0.03 $\pm$ 0.07 / 0.13 $\pm$ 0.09 0.16 $\pm$ 0.10 0.80 $\pm$ 0.18 / 0.81 $\pm$ 0.19 / 0.80 $\pm$ 0.19
GPT3.5 Turbo 0.20 $\pm$ 0.14 / 0.06 $\pm$ 0.12 / 0.18 $\pm$ 0.14 0.20 $\pm$ 0.15 0.78 $\pm$ 0.25 / 0.79 $\pm$ 0.25 / 0.78 $\pm$ 0.25 0.17 $\pm$ 0.10 / 0.04 $\pm$ 0.08 / 0.15 $\pm$ 0.10 0.18 $\pm$ 0.12 0.79 $\pm$ 0.21 / 0.80 $\pm$ 0.22 / 0.80 $\pm$ 0.21
GPT4o Mini 0.19 $\pm$ 0.11 / 0.06 $\pm$ 0.08 / 0.17 $\pm$ 0.10 0.22 $\pm$ 0.15 0.78 $\pm$ 0.24 / 0.81 $\pm$ 0.25 / 0.80 $\pm$ 0.24 0.20 $\pm$ 0.09 / 0.05 $\pm$ 0.07 / 0.17 $\pm$ 0.08 0.22 $\pm$ 0.12 0.85 $\pm$ 0.05 / 0.87 $\pm$ 0.05 / 0.86 $\pm$ 0.05
GPT4 Turbo 0.23 $\pm$ 0.13 / 0.08 $\pm$ 0.11 / 0.20 $\pm$ 0.13 0.23 $\pm$ 0.15 0.84 $\pm$ 0.16 / 0.84 $\pm$ 0.16 / 0.84 $\pm$ 0.16 0.22 $\pm$ 0.11 / 0.07 $\pm$ 0.09 / 0.19 $\pm$ 0.10 0.23 $\pm$ 0.14 0.85 $\pm$ 0.10 / 0.86 $\pm$ 0.10 / 0.86 $\pm$ 0.11
Table 14: Diversity evaluation of LLMs using Dist-1 and Dist-2 metrics across benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2 Dist-1 Dist-2
Mistral 7B 0.86 $\pm$ 0.2 0.95 $\pm$ 0.21 0.87 $\pm$ 0.2 0.95 $\pm$ 0.21 0.87 $\pm$ 0.15 0.94 $\pm$ 0.22 0.86 $\pm$ 0.2 0.95 $\pm$ 0.21 0.49 $\pm$ 0.44 0.56 $\pm$ 0.49 0.75 $\pm$ 0.26 0.88 $\pm$ 0.3
Qwen2 7B 0.90 $\pm$ 0.14 0.98 $\pm$ 0.14 0.90 $\pm$ 0.1 0.99 $\pm$ 0.09 0.90 $\pm$ 0.15 0.97 $\pm$ 0.15 0.91 $\pm$ 0.07 0.99 $\pm$ 0.06 0.43 $\pm$ 0.44 0.49 $\pm$ 0.5 0.60 $\pm$ 0.37 0.72 $\pm$ 0.44
Llama3.1 8B 0.78 $\pm$ 0.27 0.89 $\pm$ 0.3 0.89 $\pm$ 0.1 0.99 $\pm$ 0.08 0.84 $\pm$ 0.19 0.95 $\pm$ 0.2 0.90 $\pm$ 0.07 0.99 $\pm$ 0.04 0.44 $\pm$ 0.41 0.53 $\pm$ 0.48 0.76 $\pm$ 0.19 0.93 $\pm$ 0.21
GPT3.5 Turbo 0.91 $\pm$ 0.06 1.0 $\pm$ 0.01 0.91 $\pm$ 0.06 0.99 $\pm$ 0.03 0.92 $\pm$ 0.06 1.0 $\pm$ 0.01 0.87 $\pm$ 0.15 0.97 $\pm$ 0.15 0.79 $\pm$ 0.26 0.90 $\pm$ 0.28 0.78 $\pm$ 0.22 0.92 $\pm$ 0.25
GPT4o Mini 0.89 $\pm$ 0.07 0.99 $\pm$ 0.05 0.88 $\pm$ 0.06 0.99 $\pm$ 0.04 0.91 $\pm$ 0.06 1.0 $\pm$ 0.03 0.89 $\pm$ 0.05 1.0 $\pm$ 0.01 0.79 $\pm$ 0.25 0.91 $\pm$ 0.28 0.85 $\pm$ 0.07 0.99 $\pm$ 0.06
GPT4 Turbo 0.93 $\pm$ 0.05 1.0 $\pm$ 0.01 0.93 $\pm$ 0.05 1.0 $\pm$ 0.01 0.92 $\pm$ 0.05 1.0 $\pm$ 0.01 0.91 $\pm$ 0.06 1.0 $\pm$ 0.03 0.86 $\pm$ 0.18 0.96 $\pm$ 0.18 0.86 $\pm$ 0.11 0.98 $\pm$ 0.11
Each of these metrics offers distinct insights into the model’s performance, collectively ensuring a comprehensive and nuanced evaluation of the generated responses. When the model’s output couldn’t be evaluated because of parsing errors, we assigned default worst-case values. While discarding unparsable outputs, it is essential to ensure the automation of our benchmark.
7.3 Comprenehsive Results
Tables 11-17 present the detailed experimental results for all evaluated aspects, including both the average metric values and their standard deviations.
Table 15: Coherence evaluation using UE Score on benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Mistral 7B 0.46 ± 0.72 0.31 ± 0.67 0.48 ± 0.79 0.44 ± 0.79 0.21 ± 0.52 0.33 ± 0.62
Qwen2 7B 0.42 ± 0.73 0.38 ± 0.72 0.39 ± 0.75 0.52 ± 0.84 0.17 ± 0.48 0.36 ± 0.70
Llama3.1 8B 0.23 ± 0.58 0.22 ± 0.58 0.33 ± 0.70 0.37 ± 0.73 0.17 ± 0.48 0.33 ± 0.64
GPT 3.5 Turbo 0.49 ± 0.76 0.38 ± 0.74 0.51 ± 0.82 0.57 ± 0.86 0.37 ± 0.67 0.45 ± 0.72
GPT 4o Mini 0.40 ± 0.75 0.38 ± 0.74 0.38 ± 0.74 0.39 ± 0.76 0.35 ± 0.69 0.37 ± 0.70
GPT4 Turbo 0.43 ± 0.78 0.42 ± 0.78 0.56 ± 0.84 0.50 ± 0.82 0.39 ± 0.73 0.45 ± 0.76
Table 16: Personalization evaluation of LLMs using P-Dist Score and C Score across the benchmarking dataset, in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD
Mistral-7B 0.62 ±0.20 0.33 ±0.73 0.54 ±0.20 0.02 ±0.67 0.46 ±0.18 -0.00 ±0.59 0.43 ±0.17 -0.08 ±0.59 0.30 ±0.28 -0.23 ±0.82 0.53 ±0.22 0.41 ±0.75
Qwen2-7B 0.61 ±0.17 0.21 ±0.70 0.63 ±0.13 0.23 ±0.67 0.46 ±0.15 -0.11 ±0.61 0.49 ±0.13 0.01 ±0.57 0.29 ±0.30 -0.31 ±0.84 0.43 ±0.28 0.21 ±0.89
Llama3.1-8B 0.55 ±0.22 0.03 ±0.72 0.58 ±0.13 0.05 ±0.63 0.45 ±0.17 -0.11 ±0.61 0.46 ±0.14 -0.05 ±0.53 0.33 ±0.33 -0.22 ±0.87 0.59 ±0.19 0.54 ±0.68
GPT3.5-Turbo 0.63 ±0.15 0.23 ±0.69 0.60 ±0.13 0.10 ±0.62 0.47 ±0.15 -0.01 ±0.60 0.48 ±0.16 0.02 ±0.60 0.51 ±0.21 0.26 ±0.75 0.52 ±0.18 0.40 ±0.72
GPT-40-Mini 0.63 ±0.12 0.10 ±0.63 0.64 ±0.11 0.14 ±0.61 0.48 ±0.13 -0.10 ±0.52 0.50 ±0.13 -0.05 ±0.52 0.54 ±0.19 0.16 ±0.71 0.57 ±0.11 0.30 ±0.63
GPT-4-Turbo 0.60 ±0.13 0.08 ±0.62 0.59 ±0.12 0.01 ±0.58 0.47 ±0.14 -0.05 ±0.53 0.47 ±0.13 -0.05 ±0.53 0.51 ±0.15 0.14 ±0.66 0.52 ±0.13 0.21 ±0.63
Table 17: Response Time and Failure Ratio of LLMs on IT-ConvAI2, Blended Skill Talk, and FoCus datasets, in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio
Mistral 7B 3.11 ± 1.11 0.05 ± 0.00 6.90 ± 1.61 0.05 ± 0.00 3.87 ± 0.72 0.05 ± 0.00 6.71 ± 1.85 0.05 ± 0.00 4.39 ± 0.36 0.44 ± 0.00 7.67 ± 1.27 0.10 ± 0.00
Qwen2 7B 1.97 ± 0.72 0.02 ± 0.00 4.70 ± 0.94 0.01 ± 0.00 2.21 ± 0.67 0.02 ± 0.00 4.8 ± 0.88 0.00 ± 0.00 3.74 ± 0.76 0.51 ± 0.00 7.28 ± 1.07 0.27 ± 0.00
Llama3.1 8B 4.15 ± 0.53 0.10 ± 0.00 8.17 ± 0.11 0.01 ± 0.00 4.28 ± 0.26 0.04 ± 0.00 8.69 ± 0.15 0.00 ± 0.00 4.54 ± 0.21 0.45 ± 0.00 8.93 ± 0.15 0.05 ± 0.00
GPT3.5 Turbo 1.18 ± 0.57 0.00 ± 0.00 1.40 ± 0.30 0.00 ± 0.00 0.91 ± 0.21 0.00 ± 0.00 1.56 ± 0.47 0.02 ± 0.00 1.30 ± 0.42 0.09 ± 0.00 2.16 ± 0.54 0.07 ± 0.00
GPT4o Mini 0.99 ± 0.37 0.00 ± 0.00 2.20 ± 0.67 0.00 ± 0.00 0.94 ± 0.28 0.00 ± 0.00 1.63 ± 0.42 0.00 ± 0.00 1.43 ± 0.66 0.09 ± 0.00 3.05 ± 8.65 0.00 ± 0.00
GPT4 Turbo 2.21 ± 0.71 0.00 ± 0.00 4.81 ± 1.45 0.00 ± 0.00 1.84 ± 0.5 0.00 ± 0.00 5.15 ± 1.13 0.00 ± 0.00 3.32 ± 1.24 0.04 ± 0.00 7.29 ± 2.43 0.01 ± 0.00